OpenStudy (anonymous):

Assume f(x) = -2x + 8 and g(x) = 3x. What is the value of (g o f)(3)?

OpenStudy (anonymous):

A. 23 B. 6 C. -10 D. X+8

OpenStudy (anonymous):

f(g(x))=-2(3x)+8 f(g(3))=-2(3*3)+8 f(g(3))=-18+8

OpenStudy (anonymous):

Answer was 6.. Lol

OpenStudy (anonymous):

it's not 6

OpenStudy (anonymous):

ah, nevermind g(f(3))=6

OpenStudy (anonymous):

the other way around, i read it wrong